# The emergence of sparse attention: impact of data distribution and benefits of repetition

### v1

@@ Abstract
Scaling has been central to the recent success of large language models [1, 2, 3, 4, 5], with scaling laws [6, 7] describing how increased model size, data, and training consistently improve average performance. However, beneath this macroscopic predictability, model performance on specific tasks often reveals capabilities that appear suddenly beyond critical scaling thresholds – a phenomenon known as emergence [8, 9, 10, 11]. While recent studies have begun to characterize how emergence can appear after a critical training time [12, 13, 14, 15, 16, 17, 18], a comprehensive scientific understanding remains elusive.
This work explores sparse attention as a lens to understand emergence during training. The formation of sparse attention – where Transformers’ attention layers focus on a small subset of critical tokens – coincides with sudden performance improvements in many emergent behaviors, including in-context learning abilities [12, 15] and factual recall [18]. We investigate why the development of sparse attention can lead to abrupt performance improvements and reveal how characteristics of the training data influence the speed of emergence. We make two key contributions:
@@ 1 Motivation
Can emergence be predicted? Emergence in large language models not only challenges our scientific understanding of how they acquire new skills but also poses AI safety issues [19] due to its unpredictable nature, while additionally complicating frontier model development. To address this unpredictability, researchers have proposed several progress metrics under which scaling has more predictable consequences, including validation loss [20], metrics that reward partial progress [21, 22], and those that employ mechanistic interpretability-motivated measures [23]. However, a significant limitation is that these metrics typically can only be derived post-emergence [24]. An alternative approach demonstrates that fine-tuned models’ performance on certain tasks can predict whether the ability to solve the task will emerge in larger models [17]. Our work explores predictability from a different angle: understanding how the data distribution influences the learning time at which emergence occurs.
@@
Is there a link between emergence and sparse attention? The driving hypothesis underlying this work is that learning of sparse attention patterns is particularly prone to produce sharp transitions in behavior during training – i.e., the sudden emergence of new capabilities. There is a statistical argument behind this intuition: when the target a Transformer has to predict depends only on a few tokens within the context, these tokens initially have very low weight in the prediction of the model, as attention typically starts uniform. Initial progress is therefore slow and the more targeted (thus sparse) attention is, the faster learning becomes. Multiple empirical observations strengthen our hypothesis. The induction head [12], whose formation coincides with the sudden emergence of certain in-context learning abilities, fundamentally relies on the combination of two attention layers with sparse patterns. Similarly, the mechanisms underlying factual recall in large language models [23, 25] demonstrate sparse attention properties and show emergent learning dynamics [18]. These specific emergent phenomena appear to be linked to the development of sparse attention mechanisms, suggesting the existence of a potential causal relationship between the two.
@@
The intriguing interplay between repetition and emergence. While data diversity is generally considered a gold standard in machine learning, a growing body of evidence suggests that repetition can actually accelerate emergence over training. Chan et al. [13] demonstrated that showing some tasks more often favors the emergence of in-context learning. Charton and Kempe [16] revealed that repeating a subset of examples more frequently than others dramatically accelerates Transformers’ ability to solve certain arithmetic tasks. This pattern extends beyond mathematical reasoning, as Zucchet et al. [18] (and to some extent Allen-Zhu and Li [26]) showed that repeating biographies of specific individuals speeds up the development of circuits critical for factual recall from model parameters. Collectively, these results establish repetition as a fundamental property of data distributions that can systematically influence emergence timing, thus justifying integrating it in our model to better understand its role.
@@ 2.1 The single-location linear regression task
Here, the target weight matrix $W^{*}\in\mathbb{R}^{d\times d}$ is a predetermined matrix and $y\in\mathbb{R}^{d}$. To successfully solve this task, an attention-based model must learn to attend to the last token only and learn the ground-truth target weights $W^{*}$. We deliberately incorporate a sparse attention target mechanism to study the relationship between sparse attention and emergence. This task shares similarities with the single-location regression task of Marion et al. [27], although in our case, the relevant token location is always the same. The left panel of Figure 1 summarizes this task.
Figure 1: A simple task to study the emergence of sparse attention. (left) We introduce a variant of linear regression task that is analytically tractable and in which Transformers-like models need to learn sparse attention. The model must identify which token (here the last one, $x_{T}$) is relevant for the target output $y^{*}$. We incorporate two realistic forms of repetition in the data: in-context repetition, where the relevant token appears multiple times within the context, and cross-sample repetition, where an input sequence contains a special token $\tilde{x}$ (here colored in green) at the relevant position with probability $p$. See Section 2.1 for details. (right) a. As desired, the learning dynamics of our simplified Transformer (Eq. 2) exhibit a multi-phase behavior including an initial plateau, on the task without repetition ($T=512$). b. Mechanistically, the weights $w$ begin learning before attention to the relevant token $\alpha$ ($T=512$, $d=64$). Dashed lines represent optimal values. c. The duration of the initial plateau increases as a function of sequence length $T$ and input/output dimension $d$, closely following a power law scaling relationship ($R^{2}=0.999$). See Section 2.2 for details.
@@
  • In-context repetition. This occurs when specific token groups (e.g., a person’s name) repeatedly appear within a single context window. In our framework, we model this property by repeating the relevant token $x_{T}$ multiple times in the sequence $(x_{t})_{t=1}^{T}$, always at the same positions for our simplified model to be able to use it. Following Chan et al. [13], who termed this property “burstiness”, we denote $B$ as the number of times the task-relevant token appears in the context.
@@
with $a\in\mathbb{R}^{T}$ the attention scores vector and $W\in\mathbb{R}^{d\times d}$ the weight matrix. Unlike in standard Transformer attention [28], our model does not use any semantic information to determine where to attend. This simplification facilitates theoretical analysis and implicitly assumes that the attention layer has already learned to filter irrelevant contextual information. The model’s parameters $(a,W)$ are learned by minimizing the expected mean square error between predictions $y$ and targets $y^{*}$.
@@ 2.2 The learning dynamics of the simplified Transformer exhibit emerging abilities
[…]
@@
with $\alpha:=(1+(T-1)\exp(-\Delta a))^{-1}$ the attention given to the final token. In these two equations, the first term is the relevant signal and the second term is the noise coming from non-relevant tokens.
These equations enable us to elucidate the roots of emergence in this task. Initial learning is slow, as $\Delta a$ does not receive any teaching signal as $w=0$ and $w$ slowly increases as attention is initially uniform ($\alpha=1/T$). As a consequence, the feedforward weight $W$ must first align with the target weights $W^{*}$ before attention can learn. A positive feedback loop is then progressively established: increased attention improves the learning signal for $w$, and a better-learned $w$ further reinforces the correct attention pattern. This dynamic, similar to the one found in deep linear networks [29], leads to a sharp decrease in loss and the sudden emergence we observe in Figure 1.a.
@@ 2.3 Repetition speeds up emergence
[…]
@@
The role of in-context repetition is rather simple. At initialization, the attention layer allocates $B$ times more weight to the relevant tokens compared to the case without any repetition. This accelerates both initial learning of the weights and the increase of attention to the relevant tokens. Theoretically, we can show, cf. Appendix A.3, that the escape time from the initial loss plateau becomes proportional to $T\sqrt{d}/B$. Replacing $T$ by $T/B$ in the scaling law we obtained in Figure 1.c yields an almost perfect empirical fit, cf. Figure 2.b. This result highlights that $B$ reduces the sparsity of the target attention pattern and thus accelerates emergence.
@@
The role of cross-sample repetition is more intricate. Repeating one token more frequently, that is increasing $p$, causes the input covariance matrix of the relevant token, $\mathbb{E}[x_{T}x_{T}^{\top}]$, to become anisotropic. The different components of the weight matrix $W$ are then learned at different speeds. While this difference in learning speed traditionally leads to slower convergence rates in vanilla linear regression [30], it turns out to be beneficial in our attention-based version of the task. Indeed, following similar principles to the ones detailed in the previous section, learning weights in the repeated direction will lead to the attention to the relevant tokens increasing faster, and this then speeds up the learning of the non-repeated dimensions. As a consequence, the model escapes the initial learning plateau faster. Importantly, this also holds when measuring the model’s performance on non-repeated data, as seen in Figure 2. The effect of cross-sample repetition can also be understood as increasing the signal-to-noise ratio. Repeating the same token increases the amount of signal coming from the relevant token while keeping the amount of noise received from other tokens fixed, at the price of losing some information about the relevant signal.
@@
However, cross-sample repetition is not a free lunch: it only provides a temporary advantage. First, these dynamics only have a smaller test loss in the medium term. In the long run, learning is slower for similar reasons as for the standard linear regression. Second, there is an additional overfitting problem. While it does not appear in this simplified setting as anisotropic input covariance does not bias the final solution, it starts appearing for more realistic architectures. This phenomenon has been observed in practice: examples include [31, 16, 18] for synthetic tasks that have properties similar to the one we focus on here, and [32, 33] for the pretraining of large language models.
@@ 2.4 The theory qualitatively captures learning dynamics of full-fledged Transformers
We conclude our linear regression analysis by examining how our theoretical predictions extend to more realistic task versions, optimizers, and models – particularly those with standard attention that varies with inputs. Our findings indicate that learning dynamics behave qualitatively similarly, showing sharp phase transitions, with emergence time maintaining similar dependencies on data properties. However, the precise dependencies differ, which we investigate in more depth.
[…]
For this investigation, we train a standard $2$-layer $4$-heads Transformer with the Adam optimizer [34], using a constant learning rate of $10^{-4}$. Our only deviation from standard architectures is removing layer normalization, which makes solving the task more challenging. To enhance task realism, we randomly sample the positions of relevant tokens and incorporate an additional feature into the input to indicate whether a token is task-relevant. Further experimental details are provided in the appendix.
@@
For scenarios with no repetition or with in-context repetition, power laws still accurately capture the dependency of emergence time on sequence length $T$, dimension $d$ and burstiness $B$, though with significantly different exponents. Ablations (see Appendix B.3) reveal that optimizer, architecture, and task specifics all influence these exponents. Notably, replacing Adam with SGD substantially slows emergence, both in absolute terms and by increasing dependency on task difficulty. This finding illustrates a broader observation: Adam is crucial for efficient Transformer learning [e.g., 35].
@@ 3 Emergence of in-context learning, through the lens of sparse attention
We conclude by investigating to what extent the insights developed on our simple linear regression task extend to more realistic learning problems, in particular one in which in-context learning emerges. To this end, we examine an in-context associative recall task, which can be solved by learning an induction head – a well-studied circuit strongly implicated in in-context learning [12, 15], which necessitates at least two attention layers with sparse attention patterns. Our theory qualitatively captures both the learning dynamics and the impact of the training distribution on learning speed.
@@ 3.1 The in-context associative recall task
The in-context associative recall task serves as a standard benchmark for testing language models’ ability to access information in their context and to perform a simple form of one-shot learning. This task requires abilities that correlate with models’ capacity for in-context learning [12], and variations of it have been extensively studied to better understand how in-context learning emerges [13, 15, 36, 37, 38, 39, 40]. It has also been used as a testbed for recently proposed linear recurrent architectures [e.g. 41, 42, 43], as it acts as an important differentiator between attention-based and recurrent architectures [44].
@@
The number of pairs $N_{\mathrm{pairs}}$ controls the sequence length $T$ as $T=2N_{\mathrm{pairs}}+1$ (accounting for the query token), and the vocabulary size $N_{\mathrm{tokens}}$ plays a role comparable to the input dimension $d$ in the linear regression task. To model in-context repetition, we ensure that the query token appears on average $B$ times in the context (as a key). To model cross-sample repetition, we select a subset of $2$ tokens that will appear more often2222222 is an arbitrary choice that we have not ablated. and vary $p$, the probability to sample the query from this subset. The precise description of the task is provided in the appendix. The results we report in the main text are testing the learned models on data without any repetition, thereby testing the generalization abilities of models learned on repeated data.
Transformers typically solve this type of task by implementing an induction head – a circuit that combines an attention layer responsible for concatenating the representation of the current token with the previous token, and a selection attention layer responsible for retrieving the relevant information from the context. Both attention layers focus on a sparse subset of tokens (usually just one), making this task particularly well-suited to test our theory. Unlike the linear regression task where sparse attention was hard-coded in the target input-output mapping, there is no inherent constraint forcing models to develop sparse attention here. This task thus serves as the perfect testbed to evaluate our theory on more realistic, yet still finely controlled, learning scenarios.

### v2

@@ Abstract
Scaling has been central to the recent success of large language models [bengio_neural_2003, radford_language_2019, brown_language_2020, openai_gpt-4_2023, team_gemini_2023], with scaling laws [kaplan_scaling_2020, hoffmann_training_2022] describing how increased model size, data, and training consistently improve average performance. However, beneath this macroscopic predictability, model performance on specific tasks often reveals capabilities that appear suddenly beyond critical scaling thresholds – a phenomenon known as emergence [wei_emergent_2022, ganguli_predictability_2022, srivastava_beyond_2022, anderson_more_1972]. While recent studies have begun to characterize how emergence can appear after a critical training time [olsson_-context_2022, chan_data_2022, nanda_progress_2023, singh_what_2024, charton_emergent_2024, snell_predicting_2024, zucchet_how_2025], a comprehensive scientific understanding remains elusive.
This work explores sparse attention as a lens to understand emergence during training. The formation of sparse attention – where Transformers’ attention layers focus on a small subset of critical tokens – coincides with sudden performance improvements in many emergent behaviors, including in-context learning abilities [olsson_-context_2022, singh_what_2024] and factual recall [zucchet_how_2025]. We investigate why the development of sparse attention can lead to abrupt performance improvements and reveal how characteristics of the training data influence the speed of emergence. We make two key contributions:
@@ 1 Motivation
Can emergence be predicted? Emergence in large language models not only challenges our scientific understanding of how they acquire new skills but also poses AI safety issues [anwar_foundational_2024] due to its unpredictable nature, while additionally complicating frontier model development. To address this unpredictability, researchers have proposed several progress metrics under which scaling has more predictable consequences, including validation loss [du_understanding_2024], metrics that reward partial progress [schaeffer_are_2023, zhao_distributional_2025], and those that employ mechanistic interpretability-motivated measures [nanda_fact_2023]. However, a significant limitation is that these metrics typically can only be derived post-emergence [barak_emergent_2023]. An alternative approach demonstrates that fine-tuned models’ performance on certain tasks can predict whether the ability to solve the task will emerge in larger models [snell_predicting_2024]. Our work explores predictability from a different angle: understanding how the data distribution influences the learning time at which emergence occurs.
@@
Is there a link between emergence and sparse attention? The driving hypothesis underlying this work is that learning of sparse attention patterns is particularly prone to produce sharp transitions in behavior during training – i.e., the sudden emergence of new capabilities. There is a statistical argument behind this intuition: when the target a Transformer has to predict depends only on a few tokens within the context, these tokens initially have very low weight in the prediction of the model, as attention typically starts uniform. Initial progress is therefore slow and the more targeted (thus sparse) attention is, the faster learning becomes. Multiple empirical observations strengthen our hypothesis. The induction head [olsson_-context_2022], whose formation coincides with the sudden emergence of certain in-context learning abilities, fundamentally relies on the combination of two attention layers with sparse patterns. Similarly, the mechanisms underlying factual recall in large language models [nanda_fact_2023, geva_dissecting_2023] demonstrate sparse attention properties and show emergent learning dynamics [zucchet_how_2025]. These specific emergent phenomena appear to be linked to the development of sparse attention mechanisms, suggesting the existence of a potential causal relationship between the two.
@@
The intriguing interplay between repetition and emergence. While data diversity is generally considered a gold standard in machine learning, a growing body of evidence suggests that repetition can actually accelerate emergence over training. chan_data_2022 demonstrated that showing some tasks more often favors the emergence of in-context learning. charton_emergent_2024 revealed that repeating a subset of examples more frequently than others dramatically accelerates Transformers’ ability to solve certain arithmetic tasks. This pattern extends beyond mathematical reasoning, as zucchet_how_2025 (and to some extent allen-zhu_physics_2023) showed that repeating biographies of specific individuals speeds up the development of circuits critical for factual recall from model parameters. Collectively, these results establish repetition as a fundamental property of data distributions that can systematically influence emergence timing, thus justifying integrating it in our model to better understand its role.
@@ 2.1 The single-location linear regression task
Here, the target weight matrix $W^{*}\in\mathbb{R}^{d\times d}$ is a predetermined matrix and $y\in\mathbb{R}^{d}$. To successfully solve this task, an attention-based model must learn to attend to the last token only and learn the ground-truth target weights $W^{*}$. We deliberately incorporate a sparse attention target mechanism to study the relationship between sparse attention and emergence. This task shares similarities with the single-location regression task of marion_attention_2024, although in our case, the relevant token location is always the same. The left panel of Figure 1 summarizes this task.
Figure 1: A simple task to study the emergence of sparse attention. (left) We introduce a variant of linear regression task that is analytically tractable and in which Transformers-like models need to learn sparse attention. The model must identify which token (here the last one, $x_{T}$) is relevant for the target output $y^{*}$. We incorporate two realistic forms of repetition in the data: in-context repetition, where the relevant token appears multiple times within the context, and cross-sample repetition, where an input sequence contains a special token $\tilde{x}$ (here colored in green) at the relevant position with probability pp. See Section 2.1 for details. (right) a. As desired, the reduced learning dynamics of our simplified Transformer (Eq. 2) exhibit a multi-phase behavior including an initial plateau, on the task without repetition ($T=512$). b. Mechanistically, the weights $w$ begin learning before attention to the relevant token $\alpha$ ($T=512$, $d=64$). Dashed lines represent optimal values. c. The duration of the initial plateau increases as a function of sequence length $T$ and input/output dimension $d$, closely following a power law scaling relationship ($R^{2}=0.999$) that can be accurately predicted by linearizing the dynamics around initialization (Equation 14). See Section 2.2 for details.
@@
  • In-context repetition. This occurs when specific token groups (e.g., a person’s name) repeatedly appear within a single context window. In our framework, we model this property by repeating the relevant token $x_{T}$ multiple times in the sequence $(x_{t})_{t=1}^{T}$, always at the same positions for our simplified model to be able to use it. Following chan_data_2022, who termed this property “burstiness”, we denote $B$ as the number of times the task-relevant token appears in the context.
@@
with $a\in\mathbb{R}^{T}$ the attention scores vector and $W\in\mathbb{R}^{d\times d}$ the weight matrix. Unlike in standard Transformer attention [vaswani_attention_2017], our model does not use any semantic information to determine where to attend. This simplification facilitates theoretical analysis and implicitly assumes that the attention layer has already learned to filter irrelevant contextual information. The model’s parameters $(a,W)$ are learned by minimizing the expected mean square error between predictions $y$ and targets $y^{*}$.
@@ 2.2 The learning dynamics of the simplified Transformer exhibit emerging abilities
We analyze the gradient flow dynamics of the simplified model. The assumptions and mathematical details of this analysis can be found in Appendix A. Under these mild assumptions, the learning dynamics of the entire model reduces to two key scalar variables: $\Delta a:=a_{T}-a_{t}$ for any $t<T$ (all attention scores to non-relevant tokens stay the same under our assumptions) and $w$ the scalar projection111ww is formally defined as $w\coloneqq\langle W^{*},W\rangle_{F}/\lVert W^{*}\rVert_{F}\in\mathbb{R}$ with $\langle\,\cdot\,,\,\cdot\,\rangle_{F}$ the Froebenius inner product. of $W$ on $W^{*}$. These variables are initially both equal to 0 and they evolve according to the following system of ordinary differential equations:
@@
with $\alpha:=(1+(T-1)\exp(-\Delta a))^{-1}$ the attention given to the final token. In these two equations, the first term is the signal coming from the relevant token at position $T$, and the second term is the noise coming from the T1T-1 non-relevant tokens. This decomposition follows from the derivation in Appendix A.3, which groups token contributions by their relevance to the target prediction.
These equations enable us to elucidate the roots of emergence in this task. Initial learning is slow, as $\Delta a$ does not receive any teaching signal as $w=0$ and $w$ slowly increases as attention is initially uniform ($\alpha=1/T$). As a consequence, the feedforward weight $W$ must first align with the target weights $W^{*}$ before attention can learn. A positive feedback loop is then progressively established: increased attention improves the learning signal for $w$, and a better-learned $w$ further reinforces the correct attention pattern. This dynamic, similar to the one found in deep linear networks [saxe_exact_2014], leads to a sharp decrease in loss and the sudden emergence we observe in Figure 1.a.
@@ 2.3 Repetition speeds up emergence
Figure 2: Repetition speeds up emergence in the linear regression task in a theoretically predictable way. (left) Increasing in-context repetition through $B$ reduces the initial plateau, and the length of the plateau is well captured by the power law $T_{\mathrm{plateau}}=1.51\,T^{\,0.99}\,B^{-0.99}\,d^{\,0.49}$ ($R^{2}=0.999$). (right) Cross-sample repetition, modulated by the repetition probability $p$, exhibits similar effects, even when evaluating the model on a test loss without repetition (i.e., $p=0$). The length of the plateau follows $T_{\mathrm{plateau}}=2.15(\sqrt{d}T/\sqrt{p^{2}d+(1-p)^{2}})^{1.02}$ ($R^{2}=0.992$). The plateau length in both cases closely follows theoretical predictions. See Section 2.3 for details.
@@
The role of in-context repetition is rather simple. When relevant information repeats multiple times within the context, it becomes more correlated with the token to predict. This can be understood as increasing the signal to noise ratio or as effectively reducing the sequence length from $T$ to $T/B$. Since learning time scales with sequence length, this makes the task fundamentally easier. Theoretically, we can show, cf. Appendix A.3, that this intuition precisely holds and that the escape time from the initial loss plateau becomes proportional to $T\sqrt{d}/B$. Replacing $T$ by $T/B$ in the scaling law we obtained in Figure 1.c yields an almost perfect empirical fit, cf. Figure 2.b. This result highlights that $B$ reduces the sparsity of the target attention pattern and thus accelerates emergence.
@@
The role of cross-sample repetition is more intricate. Repeating one token more frequently, that is increasing $p$, causes the input covariance matrix of the relevant token, $\mathbb{E}[x_{T}x_{T}^{\top}]$, to become anisotropic. The different components of the weight matrix $W$ are then learned at different speeds. While this difference in learning speed traditionally leads to slower convergence rates in vanilla linear regression [bach_learning_2024], it turns out to be beneficial in our attention-based version of the task. Indeed, following similar principles to the ones detailed in the previous section, learning weights in the repeated direction will lead to the attention to the relevant tokens increasing faster, and this then speeds up the learning of the non-repeated dimensions. As a consequence, the model escapes the initial learning plateau faster. Importantly, this also holds when measuring the model’s performance on non-repeated data, as seen in Figure 2. The effect of cross-sample repetition can also be understood as increasing the signal-to-noise ratio. Repeating the same token increases the amount of signal coming from the relevant token while keeping the amount of noise received from other tokens fixed, at the price of losing some information about the relevant signal.
@@
However, cross-sample repetition is not a free lunch: it only provides a temporary advantage. First, these dynamics only have a smaller test loss in the medium term. In the long run, learning is slower for similar reasons as for the standard linear regression. Second, there is an additional overfitting problem. While it does not appear in this simplified setting as anisotropic input covariance does not bias the final solution, it starts appearing for more realistic architectures. This phenomenon has been observed in practice: examples include [park_competition_2024, charton_emergent_2024, zucchet_how_2025] for synthetic tasks that have properties similar to the one we focus on here, and [lee_deduplicating_2021, hernandez_scaling_2022] for the pretraining of large language models.
@@ 2.4 The theory qualitatively captures learning dynamics of full-fledged Transformers
We conclude our linear regression analysis by examining how our theoretical predictions extend to more realistic task versions, optimizers, and models – particularly those with standard attention that varies with inputs. Our findings, detailed in the remainder of this section, indicate that learning dynamics behave qualitatively similarly, showing sharp phase transitions, with emergence time maintaining similar dependencies on data properties. However, the precise dependencies differ, which we investigate in more depth.
Figure 3: Theoretical insights on the linear regression task transfer to more realistic versions of the task, the model, and the optimizer. Transformer-based architectures exhibit similar phase transitions as our toy model and its corresponding plateau length follows similar trends to the ones derived in theory. (left and middle) Evolution of the plateau length as a function of $d$, when varying $T$ and $B$ (by default $T=256$ and $B=1$). The lines corresponds to the power law $T_{\mathrm{plateau}}=0.76\,d^{\,1.29}\,T^{\,0.80}\,B^{\,-0.80}$ ($R^{2}=0.995$). (right) Same plot, this time varying the cross-sample repetition probability pp. The lines correspond to the evolution of the average plateau length as $d$ increases, for the different $p$ values. See Section 2.4 for details.
For this investigation, we train a standard $2$-layer 44-heads Transformer with the Adam optimizer [kingma_adam_2015], using a constant learning rate of $10^{-4}$. Our only deviation from standard architectures is removing layer normalization, which makes solving the task more challenging. To enhance task realism, we randomly sample the positions of relevant tokens and incorporate an additional feature into the input to indicate whether a token is task-relevant. Further experimental details are provided in the appendix.
@@
For scenarios with no repetition or with in-context repetition, power laws still accurately capture the dependency of emergence time on sequence length $T$, dimension $d$ and burstiness $B$, though with significantly different exponents. Ablations (see Appendix B.3) reveal that optimizer, architecture, and task specifics all influence these exponents. Notably, replacing Adam with SGD substantially slows emergence, both in absolute terms and by increasing dependency on task difficulty. This finding illustrates a broader observation: Adam is crucial for efficient Transformer learning [e.g., zhang_why_2024].
@@ 3 Emergence of in-context learning, through the lens of sparse attention
We conclude by investigating to what extent the insights developed on our simple linear regression task extend to more realistic learning problems, in particular one in which in-context learning emerges. To this end, we examine an in-context associative recall task, which can be solved by learning an induction head – a well-studied circuit strongly implicated in in-context learning [olsson_-context_2022, singh_what_2024], which necessitates at least two attention layers with sparse attention patterns. Our theory qualitatively captures both the learning dynamics and the impact of the training distribution on learning speed.
@@ 3.1 The in-context associative recall task
The in-context associative recall task serves as a standard benchmark for testing language models’ ability to access information in their context and to perform a simple form of one-shot learning. This task requires abilities that correlate with models’ capacity for in-context learning [olsson_-context_2022], and variations of it have been extensively studied to better understand how in-context learning emerges [chan_data_2022, singh_what_2024, reddy_mechanistic_2024, bietti_birth_2023, edelman_evolution_2024, nichani_how_2024, dangelo_selective_2025]. It has also been used as a testbed for recently proposed linear recurrent architectures [e.g. de_griffin_2024, gu_mamba_2024, yang_parallelizing_2024], as it acts as an important differentiator between attention-based and recurrent architectures [arora_zoology_2023].
@@
The number of pairs $N_{\mathrm{pairs}}$ controls the sequence length $T$ as $T=2N_{\mathrm{pairs}}+1$ (accounting for the query token), and the vocabulary size $N_{\mathrm{tokens}}$ plays a role comparable to the input dimension $d$ in the linear regression task. To model in-context repetition, we ensure that the query token appears on average $B$ times in the context (as a key). To model cross-sample repetition, we select a subset of $2$ tokens that will appear more often22222 is an arbitrary choice that we have not ablated. and vary $p$, the probability to sample the query from this subset. The precise description of the task is provided in the appendix. The results we report in the main text are testing the learned models on data without any repetition, thereby testing the generalization abilities of models learned on repeated data.
Transformers typically solve this type of task by implementing an induction head – a circuit that combines an attention layer responsible for concatenating the representation of the current token with the previous token, and a selection attention layer responsible for retrieving the relevant information from the context. Both attention layers focus on a sparse subset of tokens (usually just one), making this task particularly well-suited to test our theory. Unlike the linear regression task where sparse attention was hard-coded in the target input-output mapping, there is no inherent constraint forcing models to develop sparse attention here. We note that our toy model (Equation 2) cannot directly implement an induction head, as it lacks the relative positional information for copying and semantic similarity mechanisms for token matching. This task thus serves as the perfect testbed to show that our theoretical insights extend to more realistic, yet still finely controlled, learning scenarios in which their simplifying assumptions do not hold.